Here is a list of equations in factored form and another list in vertex form. Write each of the equations in factored form next to the equivalent equation in vertex form. Either show your work to justify each choice or explain your reason for it.
y = 2(x+4)(x–6) y=2(x+1)2 – 50
y = (x-2)(x-2) y = (x-2)2
y = (x-7)(x-11) y = (x-9)2 – 4
y = 2(x–4)(x+6) y = 2(x–9)2 –8
y = 2(x–7)(x–11) y = 2(x–1)2 – 50

Answer by lwsshak3(11628)

Here is a list of equations in factored form and another list in vertex form. Write each of the equations in factored form next to the equivalent equation in vertex form. Either show your work to justify each choice or explain your reason for it.
y = 2(x+4)(x–6) y=2(x+1)^2 – 50
y = (x-2)(x-2) y = (x-2)^2
y = (x-7)(x-11) y = (x-9)^2 – 4
y = 2(x–4)(x+6) y = 2(x–9)^2 –8
y = 2(x–7)(x–11) y = 2(x–1)^2 – 50
***
Expand and complete the square for the following:
2(x+4)(x–6)=2(x^2-2x-24)=2(x^2-2x+1)-2-24=2(x-1)^2-26
(x-2)(x-2)=(x^2-4x+4)=(x^2-4x+4)-4+4=(x-2)^2
(x-7)(x-11)=(x^2-18x+77)=(x^2-18x+81)-81+77=(x-9)^2-4
2(x–4)(x+6)=2(x^2+2x-24)=2(x^2+2x+1)-2-24=2(x+1)^2-26
2(x–7)(x–11)=2(x^2-18x+77)=2(x^2-18x+81)-162+77=2(x-9)^2-85
..
as shown by above, only following 2 pairs of equations match:
y = (x-2)(x-2) y = (x-2)^2
y = (x-7)(x-11) y = (x-9)^2 – 4